Shift report 20011014 08:00-16:00

From: Apache (apache@pii3.brahms.bnl.gov)
Date: Sun Oct 14 2001 - 16:29:20 EDT

  • Next message: Apache: "Shift report 20011014 16:00-24:00"

    ------------------------------------------------------------
    Date: Sun Oct 14 16:29:20 2001, Shift: 08:00-16:00
    Supervisor: ykl, with: 
    Log book#10 pages 1 - 5
    
    Shift Summary:
    ------------------------------------------------------------
    1. Good steady beam for five hours starting from 10 AM.
    2. Data for MRS at 45 deg 4B, FS at 12/12 deg 1/5A 
          Run 5326: Trigger1 = 457K
          Run 5329: Trigger1 = 294K   
       
       Data for MRS at 45 deg 4B, FS at 4/4 deg 1/5A
          Run 5330: Trigger 1 = 29K
       
       Spectrometer is now in the configuration:
          MRS at 45 deg 4A, FS at 4/4 deg 1/5A
    
    3. No new detector problem. Tile #34 is fixed. T2 Anode 1
       should be run at the default voltage of 1280 even ef
       the current may exceed 1 micro Amp.
    ------------------------------------------------------------
    



    This archive was generated by hypermail 2b30 : Sun Oct 14 2001 - 16:30:06 EDT